OpenStudy (firejay5):

How to get a focal point and parabolic curve out of the equation: y = (x + 1)^2?

OpenStudy (tkhunny):

First, find the vertex.

OpenStudy (anonymous):

Right and you can tell the shape of the graph from the equation given

OpenStudy (anonymous):

Find Vertex by expanding out your equation and finding your A, b and c. And use the axis of symmetry -b/2a

OpenStudy (anonymous):

this will give you your x value for vertex you find y by putting it into the equation as a Function of x

OpenStudy (tkhunny):

Normally, the Vertex Form is like this: 4p(y-k) = (x-h)^2 You have: k = 0, h = -1, and 4p = 1 or p = 1/4 The vertex is (-1,0). Use 'p' to find both Directrix and Focus.

OpenStudy (firejay5):

thank you so much @tkhunny and @Emily.lynn97

OpenStudy (anonymous):

No problem, I usually find the axis of symmetry way a little easier to find vertex if you are horrible at factoring like me lol